# The Challenges of Detecting Gases in Exoplanet Atmospheres

### v1

@@ A Case Study: The MIRI/LRS Transmission Spectrum of K2-18b
Bayesian model comparison is a powerful framework for evaluating competing hypotheses, but does not by itself constitute robust detection of chemical species or atmospheric phenomena. Model preferences are always relative to the chosen reference model and are sensitive to the structure of the explored model space. Quoted significances are arbitrary, if not meaningless, without acknowledgment of their relative nature. We highlight that reported detection significances can be strongly affected by incomplete or constrained model assumptions, especially in the low-SNR regime of exoplanet spectroscopy.
As a case study of this phenomenon we find that the JWST/MIRI transmission spectrum of K2-18 b does not constrain the planet’s atmospheric properties. Physically motivated equilibrium models do not identify meaningful constraints, and flexible retrievals admit a wide range of viable chemical compositions. We expanded the model space beyond the previously considered species. Other molecules, such as propyne (C3H4), provide comparable or greater Bayesian preference for the observed features than the proposed biosignatures DMS and DMDS[23]. Propyne is photochemically plausible[43] and has been detected in hydrogen-rich solar system planet atmospheres[47]. Moreover, it exhibits absorption features in the 1–5 m range — wavelengths being probed by current and upcoming JWST observations. These additional data may help clarify whether its presence in the atmosphere of K2-18 b remains a viable possibility. Nonetheless, given that multiple models with vastly different assumptions and compositions provide a good fit to the data, we argue that the threshold for a confident detection must be higher, not based on model comparisons only, and likely not achievable with this data alone, neither for DMS, DMDS nor propyne.
@@
Interpreting exoplanet spectra remains one of the most challenging frontiers in planetary science. Atmospheric retrieval results should not be equated with chemical detections in the absence of careful context. A true understanding of atmospheric properties requires the synthesis of diverse modeling frameworks, ranging from flexible parameterized to physically consistent, chemically coupled models. As the field matures, so too must our treatment of uncertainty. Future progress will require higher signal-to-noise observations, broader wavelength coverage to break spectral degeneracies, and continued development of rigorously benchmarked modeling tools. These are not failures, but signs of a community moving towards deeper, more reliable inference. Importantly, this is not a dispute between Bayesian and frequentist statistics; rather, we argue that robust atmospheric characterization depends on contextualizing results through multiple, complementary metrics, each addressing different questions. Just as multiple modeling approaches offer distinct insights and limitations, it is the overreliance on any single metric that risks obscuring what we actually know, and how confident we can be. In the end, a more systematic, physically motivated, and careful approach — one that explicitly acknowledges the limitations of our models and the impact of what we choose to include or omit — is essential to understanding these worlds.
@@ Acknowledgments
[…]
@@
L.W. led the development of the project and co-leads this publication with M.C.N. as both contributions were fundamental for the completion of this manuscript. L.W. and M.C.N. contributed equally to the study. L.W. performed the atmospheric modeling for the 33 parameter model, and validated the 1D-RCTE and 1D-RCPE models. L.W. contributed to the manuscript, figures, and tables preparation, interpretation of observations, and statistical analysis. M.C.N. led the parametric atmospheric modeling and led the investigation of hydrocarbons as well as the production of relevant line lists and sources of opacity. M.C.N. contributed to the manuscript, figures, and tables preparation, interpretation of observations, and statistical analysis. P.M. contributed to the discussions that shaped and formed this project, contributed to the production of atmospheric models, contributed to the statistical analysis in this manuscript, and provided critical interpretation of the results. P.M. contributed to the manuscript, figures, and tables preparation. L.T. created the grids of self-consistent models, both 1D-RCTE and 1D-RCPE. L.T. contributed to the analysis of the observations and performed part of the parameter estimation (retrievals). L.S.W. contributed to the statistical analysis and the production of the atmospheric models and grid-based methodology. Y.R. performed the GP analysis of the observations. L.T., L.S.W., and Y.R. contributed to the manuscript. A.F. contributed to the analysis of the observations, text, and figures. S.M. computed the chemical equilibrium parametric model fits with PICASO and computed 1D chemical kinetic models for K2-18 b’s atmosphere using photochem. S.M. contributed to the text, figures, and tables. S.S. helped develop the scope of the project and provided feedback to the manuscript. M.R.L. provided guidance and training to the team. M.R.L. developed the 1D-RCPE and 1D-RCTE methodology and contributed to the statistical analysis. T.B. contributed to the analysis of the observations and statistical analysis. T.B. contributed to the manuscript, figures, and tables. D.Z.S. contributed to the text and provided feedback on the scope of the project. V.P. provided comments to the manuscript and provided input into the statistical analysis in the project. D.K.S. provided comments to the manuscript.

### v3

@@ A Case Study: The MIRI/LRS Transmission Spectrum of K2-18 b
Bayesian model comparison is a powerful framework for evaluating competing hypotheses, but does not by itself constitute robust detection of chemical species or atmospheric phenomena. Model preferences are always relative to a chosen reference model and are sensitive to the structure of the explored model space. Quoted significances are arbitrary, if not meaningless, without acknowledgment of their relative nature. We highlight that detection significances can be strongly affected by incomplete or constrained model assumptions, especially in the low-SNR regime. We do not suggest a specific threshold for detection, as any recommendation is meaningless without clear acknowledgment of its relative nature. However, as the degree of confidence in any claimed detection decreases — either on a signal-to-noise ratio or model comparison basis — the more important context and the a priori plausibility of the species should become.
As a case study of this phenomenon, we find that the JWST/MIRI transmission spectrum of K2-18 b does not constrain the planet’s atmospheric properties. We expanded the model space beyond the previously considered species, finding that other molecules, such as propyne (C3H4), provide comparable or greater Bayesian preference for the observed features than the proposed biosignatures DMS and DMDS[30]. Propyne is photochemically plausible[46], has been detected in hydrogen-rich solar system planet atmospheres[50], and has been suggested to be abundant on temperate sub-Neptunes, including K2-18 b[54]. However, no molecules, including propyne, are identified with high confidence. We note that recent studies of the combined NIRISS + NIRSpec + MIRI spectrum of K2-18 b[55, 56] found evidence for CH4 and CO2 (c.f., ref[43]) due to their larger wavelength coverage than MIRI only, but they did not find statistically significant indications of DMS or DMDS. These analyses did not consider C3H4, but given the signal-to-noise ratio of the present data, alongside differing interpretations of the presence and abundance of major species including CH4, CO2 and H2O[43, 57], we expect that unambiguous identification of a less prominent molecule such as C3H4 would require a larger observational investment, possibly requiring observations from future Extremely Large Telescopes[58].
@@
Interpreting exoplanet spectra remains one of the most challenging frontiers in planetary science. Atmospheric retrieval results should not be equated with chemical detections without context. A true understanding of atmospheric properties requires the synthesis of diverse modeling frameworks, from flexible parameterized to physically consistent, chemically coupled models. As the field matures, so too must our treatment of uncertainty. Future progress will require higher signal-to-noise observations, broader wavelength coverage to break spectral degeneracies, and continued development of rigorously benchmarked modeling tools. The challenges discussed here are not failures, but signs of a community moving towards deeper, more reliable inference. In the near term, for low-SNR spectra where unique chemical attribution is not possible, the most reliable approach is to report model comparisons only as measures of relative model adequacy, and to always contextualize them with additional tests; both to evaluate whether features are distinguishable from noise and to assess whether the proposed interpretation is physically and chemically plausible relative to competing hypotheses. Importantly, this is not a dispute between Bayesian and frequentist statistics; rather, we argue that robust atmospheric characterization depends on contextualizing results through multiple, complementary metrics, each addressing different questions. Just as multiple modeling approaches offer distinct insights and limitations, it is the overreliance on any single metric that risks obscuring what we actually know, and how confident we can be. In the end, a more systematic, physically motivated, and careful approach — one that explicitly acknowledges the limitations of our models and the impact of what we choose to include or omit — is essential to understanding these worlds.
@@ Acknowledgments
L.W. and M.C.N. were supported by the Heising-Simons Foundation through a 51 Pegasi b Fellowship. P.M. and M.C.N. were supported under grant JWST-AR 06347 (PI: Nixon). P.M. acknowledges that this work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. Computing support for this work came from the Lawrence Livermore National Laboratory 19th Institutional Computing Grand Challenge program (PI: McGill). The document number is LLNL-JRNL-2005017. L.W., L.T. & M.R.L. acknowledge support from NASA XRP Grant [80NSSC24K0160] (P.I. Welbanks). D.Z.S. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303553. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. S.M. is supported by the Templeton Theory-Experiment (TEX) Cross Training Fellowship from the Templeton foundation. SM also acknowledges use of the lux supercomputer at UC Santa Cruz, funded by NSF MRI grant AST 1828315. A.D.F. acknowledges funding from NASA through the NASA Hubble Fellowship grant HST-HF2-51530.001-A awarded by STScI. L.W., L.S.W., M.R.L., Y.R. acknowledge Research Computing at Arizona State University[62] for providing HPC and storage. M.C.N. acknowledges University of Maryland high-performance computing resources used to conduct research presented in this paper. The authors thank Thomas Greene, Jonathan Lunine, and Aaron Bello-Arufe for helpful comments on the manuscript, as well as the anonymous referees for their insightful feedback.
@@
L.W. led the development of the project and co-leads this publication with M.C.N. as both contributions were fundamental for the completion of this manuscript. L.W. and M.C.N. contributed equally to the study. L.W. performed the atmospheric modeling for the 33 parameter model, and validated the 1D-RCTE and 1D-RCPE models. L.W. contributed to the manuscript, figures, and tables preparation, interpretation of observations, and statistical analysis. M.C.N. led the parametric atmospheric modeling and led the investigation of hydrocarbons as well as the production of relevant line lists and sources of opacity. M.C.N. contributed to the manuscript, figures, and tables preparation, interpretation of observations, and statistical analysis. P.M. contributed to the discussions that shaped and formed this project, contributed to the production of atmospheric models, contributed to the statistical analysis in this manuscript, and provided critical interpretation of the results. P.M. contributed to the manuscript, figures, and tables preparation. L.T. created the grids of self-consistent models, both 1D-RCTE and 1D-RCPE. L.T. contributed to the analysis of the observations and performed part of the parameter estimation (retrievals). L.S.W. contributed to the statistical analysis and the production of the atmospheric models and grid-based methodology. Y.R. performed the GP analysis of the observations. L.T., L.S.W., and Y.R. contributed to the manuscript. A.D.F. contributed to the analysis of the observations, text, and figures. S.M. computed the chemical equilibrium parametric model fits with PICASO and computed 1D chemical kinetic models for K2-18 b’s atmosphere using photochem. S.M. contributed to the text, figures, and tables. S.S. helped develop the scope of the project and provided feedback to the manuscript. M.R.L. provided guidance and training to the team. M.R.L. developed the 1D-RCPE and 1D-RCTE methodology and contributed to the statistical analysis. B.B. contributed text and comments to the manuscript. T.B. contributed to the analysis of the observations and statistical analysis. T.B. contributed to the manuscript, figures, and tables. D.Z.S. contributed to the text and provided feedback on the scope of the project. V.P. provided comments to the manuscript and provided input into the statistical analysis in the project. D.K.S. provided comments to the manuscript.